OpenStudy (anonymous):

find all real values of x such that f(x)=0 f(x)=x^2-3x-28

OpenStudy (anonymous):

(x^2-7x)+(4x-28)=0 x(x-7)+4(x-7)=0 (x+4) (x-7)=0 x=-4 or x=7

OpenStudy (anonymous):

how did you get the 7

OpenStudy (anonymous):

i got to possible numbers that when i add it gives me -3 in the original equation and when i mutiply it it gives me -28 and the two possible factors are -7 and 4

OpenStudy (anonymous):

oh you factored it??? okay makes sense now thank you

OpenStudy (anonymous):

u welcome :)

OpenStudy (anonymous):

wow i get it so much more now hey can you help me with another problem

OpenStudy (anonymous):

oh am all yours

OpenStudy (anonymous):

its just the equation that changes its f(x)= 2x+6/7 (its a fraction)

OpenStudy (anonymous):

would the domain for the first one be (-infinity,-7)(4,infinity)

OpenStudy (anonymous):

2x+6/7 (14x+6/7)=0 14x+6=7 14x=7-6 14x=1 x=1/14

OpenStudy (anonymous):

u can say all real numbers where x doesa not equal 4 and the 2nd all real numbers where x does not equal 7

OpenStudy (anonymous):

so you would use the brackets?

OpenStudy (anonymous):

on what?

OpenStudy (anonymous):

on the domain

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks a lot for your time and stuff

OpenStudy (anonymous):

u welcome :)